### v1

@@ III-B3 Illustrative Example for Interpretable Feature Extraction using LLMAs
Consider a financial analyst who receives a stream of financial news and public opinion data from social media. The state in which the financial analyst wishes to estimate if the market is in an upturn or downturn. Similar to example 1, the analyst can design questions using her domain knowledge, which extracts relevant information from the text. Since the stream of data is temporal in nature, the analyst can use our model of a LLMA, to adaptively update the belief of the underlying state, and any point interpet the interminent outputs $y$ to identify trading opportunities. We consider FNSPID, a financial news dataset [19], where we analyze news pertaining to the AAPL stock. We use Llama-3, and for each news article, ask 1616 1616 binary questions. The questions are provided in the appendix, and we plot the distribution of the difference of the ratio of close prices (a performance metric used to gauge the performance of a stock across days) in Figure 5. It is clear how LLM can be used as a sensor to parse textual observations and extract interpretable features.
@@ III-C Summary
In order to construct an interpretable model of a LLMA performing Bayesian inference, we use the LLM as a sensor attached to a Bayesian engine. This section gave a mathematical model for the LLMA, which is modeled as Bayesian sensors. The LLMA we consider in this paper is composed of an LLM and a Bayesian engine. We assumed that the entity interested in using the LLMA has access to the observation matrix described in this section, $D$. However, this might not be the case where the LLMA is used by a third party, and the Bayesian engine might not be explicit, but still, the entity might be interested in controlling and understanding the actions (decision) of the LLMA. We, therefore, discuss in the next section under what conditions we can reconstruct the utilities of the LLMA by probing the LLMA given black-box access. This extends the work done in explainable machine learning, where the deep neural network is modeled as a rationally inattentive Bayesian utility maximizer, and the post-training classifications are explained using the utilities obtained from the Bayesian revealed preferences framework [61, 62].
@@ IV-B Rationally Inattentive Bayesian Utility Maximizing Agent
where $\mathcal{E}$ is an entropic regularizer (e.g., mutual information or Renyi Entropy) [61]. Intuitively a higher information cost is incurred for a more accurate attention strategy, since we obtain a more accurate estimate of the state [62].
@@
We now remark on the correspondence between the model of the LLMA from (2) and the RIBUM tuple of (6). First note that even a single LLM $\mathcal{L}$ of the form 6 can be considered as a RIBUM where the $\mathcal{L}$ optimizes its self-attention matrix to optimize for picking the tokens that best predict the next token using the conditional probability distribution $\phi$. Further the LLMA can be considered as a product of three observation matrices: 1) from the state to the text observations 2) from the $\mathcal{L}$ (as described above) and 3) from the Bayesian engine, $D$. 1) is not in control of the LLMA; however, it is known to the analyst who simulates the text observations for a given state $x$. 2) comes from the pretraining of the LLM, which involves minimizing an entropic loss with respect to the self-attention mechanism. 3) comes from the LLMA  which has a pre-trained likelihood function on a suitable dataset.
@@ IV-C1 Necessary and Sufficient Conditions for Rationally Inattentive Bayesian Utility Maximizing Behaviour
$\displaystyle\begin{split}\sum_{u}&\max_{\bar{u}}\sum_{x}p_{l}(x,u)\hat{r}_{m}(x,\bar{u})-z_{l}\\ &-\left(\sum_{u}\sum_{x}p_{m}(x,u)\hat{r}_{m}(x,u)-z_{m}\right)\leq 0\ \forall l,m\in\mathcal{M}\end{split}$ (9)
@@
The feasibility check is summarized in Algorithm 1, derived from Theorem 1, any utility that satisfies the NIAS and NIAC inequalities with respect to the dataset $\mathbb{D}$ is a feasible utility. Hence, the Bayesian revealed preference returns a set-valued estimate (rather than point estimates) of the true utility. This set-valued estimate is given by $\hat{r}$ and the reconstructed information cost $\hat{Z}$ can be derived as [40],
$\displaystyle\begin{split}&\hat{Z}(\mathbb{D})=\max_{m\in\mathcal{M}}\biggl{(}z_{m}+\\ &\quad\sum_{u}\max_{\bar{u}\in\mathcal{U}}\sum_{x}p_{m}(x,u)\hat{r}_{m}(x,\bar{u})-\sum_{u}\sum_{x}p_{m}(x,u)\hat{r}_{m}(x,u)\biggr{)}\end{split}$ (10)
@@ IV-D Estimating Set-Valued and Point Estimates for the Utilities
Next, we describe the utility reconstruction method, which minimizes the $\ell_{1}$-norm of the reconstructed utility so as to obtain a sparse representation. Here, we manually set tolerances $\epsilon_{1}$ and $\epsilon_{2}$ for the margins with which each condition is satisfied. We minimize the following convex program, and we summarize the procedure in Algorithm 3,
@@ IV-F Summary
The LLM of a LLMA uses a entropic regularization method to provide its output, this motivates looking at the LLMA from the lens of rationally inattentive Bayesian utility maximization, which is a form of entropic regularized utility maximization. This section discussed the necessary and sufficient conditions for a LLMA to be a rationally inattentive Bayesian utility maximizer (RIBUM). We proposed two algorithms to reconstruct a point-valued estimate of the utility of the LLMA if it is a (RIBUM). We illustrate how the reconstructed utility is useful in analyzing the behavior of a black-box LLMA.
@@ V Bayesian Social Learning in A Sequence of LLM Agents
LLMs are already trained on synthetically generated data by other models [67] and also often use the output of other LLMs to output based on the current context [23]. Motivated by studying interacting LLMs, each of which has computational and privacy constraints, this section introduces a second layer of abstraction, wherein we study Bayesian social learning in a sequence of large language model agents. We first motivate the setting where a sequence of LLMAs sequentially estimate a state from their private observations and take a public action, which is used to update the public prior. We discuss the Bayesian social learning protocol in a sequence of LLMAs, which aim to detect an underlying state by sequentially analyzing text observations of the text. The optimal update equation for the public prior is derived. We consider two scenarios, one where no private observations are shared and one where private observations are shared to the next $M$ LLMAs. Finally, we show that under both scenarios, an information cascade takes place, and the agents take the same action irrespective of their private observation. To show this, we use the martingale convergence theorem [42]. We illustrate the effect of the number of private observations revealed and the resolution of the probe on the convergence in herding. We also present the mathematical model for incentivized autonomous LLM agents used later in Section IX, which is motivated by different entities employing such agents to perform Bayesian state estimation using textual data.
@@
Even if a single LLMA is used in an application, it can be treated as a sequence of different LLMAs since the context of the previous LLMA evaluation might not be available due to privacy of the content and finite-context length [24]. We consider two scenarios, one where no private observation is shared between the LLMAs and the second where each LLMAs can observe previous $M$ agents. This is motivated by practical constraints from the perspective of privacy, context length, and cost incurred inherent in using LLMAs.
@@ V-A2 Limited Context Length and Storage Constraints
Another constraint is the limited context length (length of text that the LLM can process at a time) that is inherent to the LLM used in the LLMA. Note that there are methods which allow for infinite context window or a very large (11 11 million) context window, however these take a lot of time to work which is often not feasible in a real-time Bayesian inference setting. Also, the quality of the responses decreases with increasing context [35]. There are storage constraints that won’t allow storage of an arbitrary amount of private observations, especially if there are methods (like those presented in this paper) that don’t require storage of the private observations.
@@ V-A4 Single LLMA can be modeled as a sequence of LLMAs
Finally, we remark that a single LLMA can be modeled as a sequence of LLMAs, especially given the above three constraints. This is because when a single LLMA is used to sequentially perform Bayesian inference, the constraints from above enforce that no more than $M$ observations can be considered at any given time. We consider updating the prior based on the action of the $M+1$-th previous LLMA. This might seem counterintuitive assumption, however in practical applications often the observations are processed in batches.
@@ V-B Social learning protocol when no private observation is shared
where $\pi_{k}(x)$ denotes the public prior belief over the state space $\mathcal{X}$. The agent $k$ takes an action $u\in\mathcal{U}$ minimizing the expected cost with respect to the posterior  [^7] A tie-breaking rule such as uniform sampling can be used if two actions have the same cost.,
@@
We relax this assumption in the next subsection by allowing the private observations to be shared with the next $M$ agents, which allows us to model more realistic Bayesian social learning in LLMAs. Based on the action $u_{k}$ of the agent $k$, the public prior belief $\pi_{k}$ is updated using the following filtering equation, which follows from the filtering equation of a hidden Markov model derived in Appendix A-B,
@@
$\displaystyle\begin{split}\mathds{P}(u|x=i,\pi)=\sum_{y\in\mathcal{Y}}\mathds{P}(u|y,\pi)\mathds{P}(y|x=i,\pi)\\ \mathds{P}(u|y,\pi)=\begin{cases}1,&\text{if }c^{\prime}_{u}B_{y}\pi\leq c^{\prime}_{\tilde{u}}B_{y}\pi,\tilde{u}\in\mathcal{U}\\ 0,&\text{otherwise}\end{cases}\\ \end{split}$ (17)
@@
7:     Agents $k+M+1,\dots$ update public prior using (15)
@@
V-C Social learning protocol when the last $M$ private observations are shared
We now consider a modification of the Bayesian social learning described in the previous subsection. We let the agents use the observations from the last $M$ LLMAs to update their posterior. We weaken assumption (B1) to the following,
@@ V-C Social learning protocol when the last $M$ private observations are shared
    The observations $z_{m}$ and $y_{m}$ are visible to agent $m$ and the next $M$ agents, i.e., to LLMAs $k=m,m+1,\dots,m+M$.
@@
Agent $k$ creates a vectors using the $M+1$ observations $\mathbf{y_{k}}=[y_{k-M},\dots,y_{k}]^{\prime}$. Since each of the observations of $\mathbf{y_{k}}$ of the state $x$ are sampled independently. The augmented observation space is $\mathcal{Y}^{M}$ The joint likelihood can be computed as $\mathds{P}(\mathbf{y_{k}}|x)=\prod_{m=0}^{M}B_{y_{k-m},x}$. Then the Bayesian update of (13) can be augmented as follows,
@@
    In lieu of observations $y_{m-M},\dots,y_{k-1}$, LLMA  $m$ disregards observed actions $u_{m-M},\dots,u_{m-1}$ of the previous $M$ agents.
Hence as a consequence of (B3), the action of agent $k$ is used by agents $k+M,k+M+1,\dots$ to update their public prior $\pi_{m},m=k+M,k+M+1,\dots$ using (15).
V-D Information Cascades and Herding Happen In Finite Time
This section proves that the LLMAs described in the previous section form an information cascade and herd in their actions when the public prior gets strong.
@@ V-D Information Cascades and Herding Happen In Finite Time
Information cascade implies that the public prior freezes after time $K$, and since the public prior freezes, the optimal action taken using (14) under any the posterior of (13). Since the information cascade implies the optimal action remains the same, the following definition naturally describes herding at time $K$ for LLMAs where the actions remain the same.
@@
Figure 8: Herding and Learning Regions for state space with 22 22 states. If the prior is in the region $\pi\in[\bar{\pi}_{1},\bar{\pi}_{2}]$, learning happens; otherwise, the LLMAs form an information cascade and herd in their actions.
@@
We first mathematically describe the different regions with respect to the public prior. Then, we derive the relation between the threshold of the public prior and the observation matrix for the different observations. Such a derivation can be used to see the effect of more accurate observations either by considering a higher resolution probe or by considering more number of observations.
We can derive the following for the different regions with respect to the public prior $\pi$ where herding happens and where it does not, i.e., where learning happens.
@@
$\displaystyle\implies\frac{(1-p)B_{y,2}-pB_{y,1}}{pB_{y,1}+(1-p)B_{y,2}}\leq 0\implies p\geq\frac{B_{y,2}}{B_{y,2}+B_{y,1}}$
@@
This discussion shows us that even for a simplistic setup, improving the probe accuracy of the LLM helps reduce the herding threshold. However, more accurate LLMs are often larger and have a higher unit cost, clearly highlighting the tradeoff between herding and the cost incurred. A similar result can be derived when the number of shared observations $M$ increases, as the observation space grows with $M$.

### v2

@@ III-B3 Illustrative Example for Interpretable Feature Extraction using LLMAs
Consider a financial analyst who receives a stream of financial news and public opinion data from social media. The state in which the financial analyst wishes to estimate if the market is in an upturn or downturn. Similar to example 1, the analyst can design questions using her domain knowledge, which extracts relevant information from the text. Since the stream of data is temporal in nature, the analyst can use our model of a LLMA, to adaptively update the belief of the underlying state, and any point interpet the interminent outputs $y$ to identify trading opportunities. We consider FNSPID, a financial news dataset [78], where we analyze news pertaining to the AAPL stock. We use LLaMA-3, and for each news article, ask 1616 1616 binary questions. The questions are provided in the appendix, and we plot the distribution of the difference of the ratio of close prices (a performance metric used to gauge the performance of a stock across days) in Figure 5. It is clear how LLM can be used as a sensor to parse textual observations and extract interpretable features. The illustrative example using the financial news dataset focuses on the AAPL stock to demonstrate the applicability of the study to readers who may be less familiar with the subject or come from a different disciplinary background, such as finance. We note that the period that we consider was during COVID-19, and therefore was an exception to the standard stock market behaviour. However we again highlight that the purpose of our methods is to enable analysis by providing an interpretable toolbox and not to analyse any specific event.
@@ III-C Summary
In order to construct an interpretable model of a LLMA performing Bayesian inference, we use the LLM as a sensor attached to a Bayesian engine. This section gave a mathematical model for the LLMA, which is modeled as Bayesian sensors. The LLMA we consider in this paper is composed of an LLM and a Bayesian engine. We assumed that the entity interested in using the LLMA has access to the observation matrix described in this section, $O$. However, this might not be the case where the LLMA is used by a third party, and the Bayesian engine might not be explicit, but still, the entity might be interested in controlling and understanding the actions (decision) of the LLMA. We, therefore, discuss in the next section under what conditions we can reconstruct the utilities of the LLMA by probing the LLMA given blackbox access. This extends the work done in explainable machine learning, where the deep neural network is modeled as a rationally inattentive Bayesian utility maximizer, and the post-training classifications are explained using the utilities obtained from the Bayesian revealed preferences framework [47, 46].
@@ IV-B Rationally Inattentive Bayesian Utility Maximizing Agent
where $\mathcal{E}$ is an entropic regularizer (e.g., mutual information or Renyi Entropy) [47]. Intuitively, a higher information cost is incurred for a more accurate attention strategy since we obtain a more accurate estimate of the state [46].
@@
We now remark on the correspondence between the model of the LLMA from (2) and the RIBUM tuple of (6). First note that even a single LLM $\mathcal{L}$ of the form 6 can be considered as a RIBUM where the $\mathcal{L}$ optimizes its self-attention matrix to optimize for picking the tokens that best predict the next token using the conditional probability distribution $\phi$. Further the LLMA can be considered as a product of three observation matrices: 1) from the state to the text observations 2) from the $\mathcal{L}$ (as described above) and 3) from the Bayesian engine, $O$. 1) is not in control of the LLMA; however, it is known to the analyst who simulates the text observations for a given state $x$. 2) comes from the pretraining of the LLM, which involves minimizing an entropic loss with respect to the self-attention mechanism. 3) comes from the LLMA  which has a pre-trained likelihood function on a suitable dataset.
@@ IV-C1 Necessary and Sufficient Conditions for Rationally Inattentive Bayesian Utility Maximizing Behaviour
$\displaystyle\begin{split}\sum_{u}&\max_{\bar{u}}\sum_{x}p_{l}(x,u)\hat{r}_{m}(x,\bar{u})-z_{l}\\ &-\left(\sum_{u}\sum_{x}p_{m}(x,u)\hat{r}_{m}(x,u)-z_{m}\right)\leq 0\ \forall l,m\in\mathcal{M}.\end{split}$ (9)
@@
The feasibility check is summarized in Algorithm 1 which is derived from Theorem 1. Any utility that satisfies the NIAS and NIAC inequalities with respect to the dataset $\mathbb{D}$ is a feasible utility. Hence, the Bayesian revealed preference returns a set-valued estimate (rather than point estimates) of the true utility. This set-valued estimate is given by $\hat{r}$ and the reconstructed information cost $\hat{Z}$ can be derived as [5],
$\displaystyle\begin{split}&\hat{Z}(\mathbb{D})=\max_{m\in\mathcal{M}}\biggl{(}z_{m}+\\ \ &\sum_{u}\max_{\bar{u}\in\mathcal{U}}\sum_{x}p_{m}(x,u)\hat{r}_{m}(x,\bar{u})-\sum_{u}\sum_{x}p_{m}(x,u)\hat{r}_{m}(x,u)\biggr{)}.\end{split}$ (10)
@@ IV-D Estimating Set-Valued and Point Estimates for the Utilities
Next, we describe the utility reconstruction method, which minimizes the $\ell_{1}$-norm of the reconstructed utility so as to obtain a sparse representation. Here, we manually set tolerances ϵ1+subscriptitalic-$\epsilon_{1}\in\mathbb{R}^{+}$ and ϵ2+subscriptitalic-$\epsilon_{2}\in\mathbb{R}^{+}$ for the margins with which each condition is satisfied. Ideally, the margin should be as high as possible such that the linear inequalities still have a feasible solution. Therefore, some trial and error is required to select a suitable $\epsilon_{1}$ and $\epsilon_{2}$. However, this is possible since the Bayesian revealed preference step is offline. We minimize the following convex program, and we summarize the procedure in Algorithm 3,
@@
Remark 16.
During pre-training, LLMs are trained on massive datasets like the CommonCrawl, which are of the order of hundreds of terabytes, and therefore, they can produce outputs for different contexts. Therefore there is no sample size restriction on the dataset for applying Bayesian revealed preferences to LLMAs. It is important to emphasize the Bayesian revealed preferences provide a necessary and sufficient condition for Bayesian utility maximization, thereby providing a rigorous data analysis tool for real-world data. The primary requirement of both algorithms is that the dataset must be collected in at least two environments, with the action posterior properly defined. There are no additional prerequisites for reconstructing the set-valued utilities from the action posteriors. However, in practice, when the action posterior is estimated empirically, errors are introduced due to finite sample approximations. Characterizing the exact convergence of empirical action posterior probabilities to their true values would be an intriguing direction for future research.
@@ IV-F Summary
The LLM of a LLMA uses a entropic regularization method to provide its output, this motivates looking at the LLMA from the lens of rationally inattentive Bayesian utility maximization, which is a form of entropic regularized utility maximization. This section discussed the necessary and sufficient conditions for a LLMA to be a rationally inattentive Bayesian utility maximizer (RIBUM). We proposed two algorithms to reconstruct a point-valued estimate of the utility of the LLMA if it is a (RIBUM). We illustrate how the reconstructed utility is useful in analyzing the behavior of a blackbox LLMA.
@@ V Bayesian Social Learning in A Sequence of LLM Agents
LLMs are already trained on synthetically generated data by other models [15] and also often use the output of other LLMs to output based on the current context [29]. Motivated by studying interacting LLMs, each of which has computational and privacy constraints, this section introduces a second layer of abstraction, wherein we study Bayesian social learning in a sequence of large language model agents. We first motivate the setting where a sequence of LLMAs sequentially estimate a state from their private observations and take a public action, which is used to update the public belief. We discuss the Bayesian social learning protocol in a sequence of LLMAs, which aim to detect an underlying state by sequentially analyzing text observations of the text. The optimal update equation for the public belief is derived. We consider two scenarios, one where no private observations are shared and one where private observations are shared to the next $L$ LLMAs. Finally, we show that under both scenarios, an information cascade takes place, and the agents take the same action irrespective of their private observation. To show this, we use the martingale convergence theorem [79]. We illustrate the effect of the number of private observations revealed and the resolution of the probe on the convergence in herding. We also present the mathematical model for incentivized autonomous LLM agents used later in Section IX, which is motivated by different entities employing such agents to perform Bayesian state estimation using textual data.
@@
Even if a single LLMA is used in an application, it can be treated as a sequence of different LLMAs since the context of the previous LLMA evaluation might not be available due to privacy of the content and finite-context length [25]. We consider two scenarios, one where no private observation is shared between the LLMAs and the second where each LLMAs can observe previous $L$ agents. This is motivated by practical constraints from the perspective of privacy, context length, and cost incurred inherent in using LLMAs.
@@ V-A2 Limited Context Length and Storage Constraints
Another constraint is the limited context length (length of text that the LLM can process at a time) that is inherent to the LLM used in the LLMA. Note that there are methods which allow for infinite context window or a very large (11 11 million) context window, however these take a lot of time to work which is often not feasible in a real-time Bayesian inference setting. Also, the quality of the responses decreases with increasing context [19]. There are storage constraints that won’t allow storage of an arbitrary amount of private observations, especially if there are methods (like those presented in this paper) that do not require storage of the private observations.
@@ V-A4 Single LLMA can be modeled as a sequence of LLMAs
Finally, we remark that a single LLMA can be modeled as a sequence of LLMAs, especially given the above three constraints. This is because when a single LLMA is used to sequentially perform Bayesian inference, the constraints from above enforce that no more than $L$ observations can be considered at any given time. We consider updating the prior based on the action of the $L+1$-th previous LLMA. This might seem counterintuitive assumption, however in practical applications often the observations are processed in batches.
@@ V-B Social learning protocol when no private observation is shared
where $\pi_{k}(x)$ denotes the public belief belief over the state space $\mathcal{X}$. The agent $k$ takes an action $u\in\mathcal{U}$ minimizing the expected cost with respect to the posterior [^8] A tie-breaking rule such as uniform sampling can be used if two actions have the same cost.,
@@
We relax this assumption in the next subsection by allowing the private observations to be shared with the next $L$ agents, which allows us to model more realistic Bayesian social learning in LLMAs. Based on the action $u_{k}$ of the agent $k$, the public belief belief $\pi_{k}$ is updated using the following filtering equation, which follows from the filtering equation of a hidden Markov model derived in Appendix A-B,
@@
$\displaystyle\begin{split}\mathds{P}(u|x=i,\pi)=\sum_{y\in\mathcal{Y}}\mathds{P}(u|y,\pi)\mathds{P}(y|x=i,\pi),\\ \mathds{P}(u|y,\pi)=\begin{cases}1,&\text{if }c^{\prime}_{u}B_{y}\pi\leq c^{\prime}_{\tilde{u}}B_{y}\pi,\tilde{u}\in\mathcal{U}\\ 0,&\text{otherwise}\end{cases},\\ \end{split}$ (17)
@@
7:     Agents $k+L+1,\dots$ update public belief using (15)
@@
V-C Social learning protocol when the last $L$ private observations are shared
We now consider a modification of the Bayesian social learning described in the previous subsection. We let the agents use the observations from the last $L$ LLMAs to update their posterior. We weaken assumption (B1) to the following,
@@ V-C Social learning protocol when the last $L$ private observations are shared
    The observations $z_{m}$ and $y_{m}$ are visible to agent $m$ and the next $L$ agents, i.e., to LLMAs $k=m,m+1,\dots,m+L$.
@@
Agent $k$ creates a vectors using the $L+1$ observations $\mathbf{y_{k}}=[y_{k-L},\dots,y_{k}]^{\prime}$. Since each of the observations of $\mathbf{y_{k}}$ of the state $x$ are sampled independently. The augmented observation space is $\mathcal{Y}^{L}$ The joint likelihood can be computed as $\mathds{P}(\mathbf{y_{k}}|x)=\prod_{m=0}^{L}B_{y_{k-m},x}$. Then the Bayesian update of (13) can be augmented as follows,
@@
    In lieu of observations $y_{m-L},\dots,y_{k-1}$, LLMA  $m$ disregards observed actions $u_{m-L},\dots,u_{m-1}$ of the previous $L$ agents.
Hence as a consequence of (B3), the action of agent $k$ is used by agents $k+L,k+L+1,\dots$ to update their public belief $\pi_{m},m=k+L,k+L+1,\dots$ using (15).
V-D Emergence of Herds and Information Cascades
This section proves that the LLMAs described in the previous section form an information cascade and herd in their actions when the public belief gets strong.
@@ V-D Emergence of Herds and Information Cascades
Information cascade implies that the public belief freezes after time $K$, and since the public belief freezes, the optimal action taken using (14) under any the posterior of (13). Since the information cascade implies the optimal action remains the same, the following definition naturally describes herding at time $K$ for LLMAs where the actions remain the same.
@@
Figure 8: Herding and Learning Regions for state space with 22 22 states. The scalar $\pi_{k}\in[0,1]$ denotes the prior of state 11 11. If the prior is in the region $\pi\in[\bar{\pi}_{1},\bar{\pi}_{2}]$, learning happens; otherwise, the LLMAs form an information cascade and herd in their actions.
@@
We first mathematically describe the different regions with respect to the public belief. Then, we derive the relation between the threshold of the public belief and the observation matrix for the different observations. Such a derivation can be used to see the effect of more accurate observations either by considering a higher resolution probe or by considering more number of observations.
We can derive the following for the different regions with respect to the public belief $\pi$ where herding happens and where it does not, i.e., where learning happens.
@@
$\displaystyle\implies\frac{(1-p)B_{y,2}-pB_{y,1}}{pB_{y,1}+(1-p)B_{y,2}}\leq 0\implies p\geq\frac{B_{y,2}}{B_{y,2}+B_{y,1}},$
@@
This discussion shows us that even for a simplistic setup, improving the probe accuracy of the LLM helps reduce the herding threshold. However, more accurate LLMs are often larger and have a higher unit cost, clearly highlighting the tradeoff between herding and the cost incurred. A similar result can be derived when the number of shared observations $L$ increases, as the observation space grows with $L$.